OpenStudy (anonymous):

Consider the function represented by 9x + 3y = 12 with x as the independent variable. How can this function be written using function notation? f(x) = ______ x + 4

OpenStudy (anonymous):

@mathstudent55 @TheSmartOne @iGreen

OpenStudy (irishboy123):

**9x + 3y = 12** once you get this looking like **y = [insert whatever your answer is here]** you are good to go all the usual rules of algebraic manipulation apply. your start point might be to divide everything by 3.

OpenStudy (anonymous):

3x?

OpenStudy (anonymous):

am i right ?

OpenStudy (irishboy123):

you are asked to produce an equation that reads: f(x) = ______ x + 4 ie f(x) = (something * x) + 4 which means y = (something * x) + 4 you know that: 9x + 3y = 12 how do you make that say: y = ........ ?

OpenStudy (anonymous):

8?

OpenStudy (irishboy123):

the answer is an equation, not a number. you know that: 9x + 3y = 12 what does y = ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

2x+4

OpenStudy (irishboy123):

nah!!! let's do this: 9x + 3y = 12 9x - 9x + 3y = 12 - 9x 3y = 12 - 9x so y = ??? and the answer is not 2 or 7 or 27 or 5 or 999 or -26. the answer involves x.

OpenStudy (anonymous):

y=4-3x

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

so f(x) = ??